40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 11 10:58:58 2022, Tega, Update, SUS, Temporary watchdog 
    Reply  Tue Jan 11 13:43:14 2022, Koji, Update, SUS, Temporary watchdog 
    Reply  Wed Jan 19 21:39:22 2022, Tega, Update, SUS, Temporary watchdog 
       Reply  Thu Jan 20 00:26:50 2022, Koji, Update, SUS, Temporary watchdog 
       Reply  Thu Jan 20 17:21:21 2022, Tega, Update, SUS, Temporary watchdog 
Message ID: 16606     Entry time: Thu Jan 20 17:21:21 2022     In reply to: 16600
Author: Tega 
Type: Update 
Category: SUS 
Subject: Temporary watchdog 

Temp software watchdog now operational for LO1 and the remaining optics!

Koji helped me understand how to write to switches and we tried for a while to only turnoff the output switch of the filters instead of the writing a zero that resets everything in the filter.

Eventually, I was able to move this effort foward by realising that I can pass the control trigger along multiple records using the forwarding option 'FLNK'. When I added this field to the trigger block, record(dfanout,"C1:SUS-LO1_PUSH_ALL"), and subsequent calculation blocks, record(calcout,"C1:SUS-LO1_COILSWa") to record(calcout,"C1:SUS-LO1_COILSWd"), everything started working right.

Quote:

After some work on the reference database file, we now have a template for temporary watchdog implementation for LO1 located here "/cvs/cds/caltech/target/c1susaux/C1_SUS-AUX_LO1.db".

Basically, what I have done is swap the EPICS asyn analog input readout for the COIL and OSEM to accessible medm channels, then write out watchdog enable/disable to coil filter SW2 switch. Everything else in the file remains the same. I am worried about some of the conversions but the only way to know more is to see the output on the medm screen.

To test, I restarted c1su2 but this did not make the LO1 database available, so I am guessing that we also need to restart the c1sus, which can be done tomorrow.

 

ELOG V3.1.3-